Express −1 + i in polar form. Hence show that (−1 + i )^16 is real and that 1/(−1 + i )^6
is purely imaginary, giving the value for each.
1
Expert's answer
2020-05-25T19:09:33-0400

"let \ z=-1+i\ , \ x=-1\ and \ y=1\\

\text{we get}\\
 \ r= \sqrt{x^2+y^2}=\sqrt{1+1}=\sqrt{2} \ and\\

\ \theta=\arctan \frac{y}{x}=\arctan \frac{1}{-1}=\frac{3 \pi}{4}\\

\text{Since, in the polar form, we have}\\

\ z=r(\cos \theta+i \sin \theta)\ .\\



\text{So, we get}\\

\ z=\sqrt{2}(\cos \frac{3 \pi}{4}+i \sin\frac{3 \pi}{4}).\\
\text{and}\\
\ z^{16}=(\sqrt{2})^{16}(\cos \frac{3 \pi}{4}+i \sin\frac{3 \pi}{4})^{16}\ .\\

\text{From De Moivre's Theorem, we have}\\

(\cos \theta+i \sin \theta)^{n}=\cos n \theta+i \sin n \theta\\

\text{So, we get}\\

\ z^{16}=(\sqrt{2})^{16}(\cos 12 \pi+i \sin 12 \pi)\\
\ \ \ \ \ \ \ =(\sqrt{2})^{16}(\cos 0+i \sin 0)=256\ .\\



\text{Also, we get}\\

\ \frac{1}{z^{6}}=z^{-6}=(\sqrt{2})^{-6}(\cos \frac{3 \pi}{4}+i \sin\frac{3 \pi}{4})^{-6}\ .\\

\ z^{-6}=(\sqrt{2})^{-6}(\cos \frac{-9 \pi}{2}+i \sin \frac{-9 \pi}{2})\\
\ \ \ \ \ \ \ \ =(\sqrt{2})^{-6}(\cos \frac{- \pi}{2}+i \sin \frac{- \pi}{2})=-\frac{1}{8}i\ .\\

\text{So, we can conclude that}\\

1. \ (-1+i)^{16}\text{ is real and equal} \ 256\ .\\

2.\ (-1+i)^{-6} \text{ is purely imaginary and equal} \ -\frac{1}{8}i\ \\"